# Nearly Minimax Optimal Regret for Multinomial Logistic Bandit

### v6

@@ F.1 Main Proof of Lemma 1
where the first inequality holds since $H_{s}\succeq\lambda\mathbf{I}_{d}$. By setting $\eta=\frac{1}{2}\log(K+1)+2$ and $\lambda=84\sqrt{2}d\eta$, we derive that
$\|\mathbf{w}_{t}-\mathbf{w}^{\star}\|_{H_{t}}\leq\beta_{t}(\delta)=\mathcal{O}\left(\sqrt{d}\log t\log K\right),$
which conclude the proof of Lemma 1. ∎
Remark F.1.
If the boundedness assumption on the parameter is relaxed to $\|\mathbf{w}\|_{2}\leq B$, then $\beta_{t}(\delta)=\mathcal{O}\left(B\sqrt{d}\log t\log K\right)$.
@@ F.2 Proofs of Lemmas for Lemma 1
Let ~$\tilde{\ell}_{s}(\mathbf{w})=\ell_{s}(\mathbf{w}_{s})+\langle\mathbf{w}-\mathbf{w}_{s},\nabla\ell_{s}(\mathbf{w}_{s})\rangle+\frac{1}{2}\|\mathbf{w}-\mathbf{w}_{s}\|^{2}_{\nabla^{2}\ell_{s}(\mathbf{w}_{s})}$ be a second-order approximation of the original function $\ell_{s}(\mathbf{w})$ at the point $\mathbf{w}_{s}$. The update rule (3) can also be expressed as
@@
~$\langle\nabla\tilde{\ell}_{s}(\mathbf{w}_{s+1}),\mathbf{w}_{s+1}-\mathbf{w}^{\star}\rangle\leq\frac{1}{2\eta}\left(\|\mathbf{w}_{s}-\mathbf{w}^{\star}\|_{H_{s}}^{2}-\|\mathbf{w}_{s+1}-\mathbf{w}^{\star}\|_{H_{s}}^{2}-\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\|_{H_{s}}^{2}\right).$ (F.4)
@@
where $\zeta=\log(K+1)+4$. Then, by combining (F.4) and (F.5), we have
@@
Then, one can easily verify that $\|\mathbf{d}^{\prime}_{s}\|_{\infty}\leq 1$ since $\|\mathbf{z}_{s}^{\star}\|_{\infty}\leq\max_{i\in S_{s}}\|x_{xi}\|_{2}\|\mathbf{w}^{\star}\|_{2}\leq 1$ and $\|\tilde{\mathbf{z}}_{s}^{\mu}\|_{\infty}\leq\log(1+(K+1)/\mu)$. On the other hand, $\mathbf{d}_{s}^{\prime}$ is $\mathcal{F}_{s}$-measurable since $\mathbf{z}_{s}^{\star}$ and $\tilde{\mathbf{z}}_{s}^{\mu}$ are independent of $\mathbf{y}_{s}$. Moreover, we have $\|\mathbf{d}^{\prime}_{s}\|^{2}_{\mathbb{E}[\bm{\varepsilon}_{s}\bm{\varepsilon}_{s}^{\top}\mid\mathcal{F}_{s}]}=\|\mathbf{d}_{s}\|^{2}_{\mathbb{E}\left[(\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s})(\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s})^{\top}\mid\mathcal{F}_{s}\right]}=\|\mathbf{d}_{s}\|^{2}_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}$ and $\|\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s}\|_{1}\leq 2$. Thus, by Lemma F.5, with probability at least $1-\delta$, for any $t\geq 1$, we have
@@
$\leq(c_{\mu}+1)\sqrt{\lambda+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}\left(\frac{\sqrt{\lambda}}{4}+\frac{4}{\sqrt{\lambda}}\log\cdot\left(\frac{2\sqrt{1+\frac{1}{\lambda}\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}}{\delta}\right)\right)$
$\leq(c_{\mu}+1)\sqrt{\lambda+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}\cdot\left(\frac{\sqrt{\lambda}}{4}+4\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right),$ (F.8)
where the second inequality holds because $\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}=\mathbf{d}_{s}^{\top}\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})\mathbf{d}_{s}\leq 2$ and $\lambda\geq 1$. Then, combining (F.8) and (F.7), and rearranging the terms, we obtain
@@
$\leq(c_{\mu}+1)\sqrt{\lambda+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}\cdot\left(\frac{\sqrt{\lambda}}{4}+4\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)-\sum_{s=1}^{t}\frac{1}{c_{\mu}}\|\mathbf{z}_{s}^{\star}-\tilde{\mathbf{z}}_{s}^{\mu}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}$
$\leq(c_{\mu}+1)\sqrt{\lambda+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}\cdot\left(\frac{\sqrt{\lambda}}{4}+4\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)-(c_{\mu}+1)\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}$
$\leq(c_{\mu}+1)\left(\lambda+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}\right)+(c_{\mu}+1)\left(\frac{\sqrt{\lambda}}{4}+4\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}$
@@
$=(c_{\mu}+1)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right),$ (F.9)
where the third inequality holds due to the AM-GM inequality. Finally, combining (F.6) and (F.9), by setting $\mu=1/t$, we have
$\sum_{s=1}^{t}\left(\ell_{s}(\mathbf{w}^{\star})-\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})\right)$
$\leq(c_{\mu}+1)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right)+2\mu t$
$\leq\left(3\log(1+(K+1)t)+3\right)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right)+2$
@@
$L_{s}(\mathbf{w})\leq\tilde{L}_{s}(\mathbf{w})+e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla\ell_{s}(\mathbf{w}_{s+1})}^{2}.$ (F.12)
We define the function $\tilde{f}_{s}:\mathcal{W}\rightarrow\mathbb{R}$ as
@@
Then, we can establish a lower bound for the expectation in (F.10) as follows:
@@
where the first inequality holds by (F.12) and the last equality holds by (F.11). We define $\tilde{Z}_{s+1}=\int_{\mathbb{R}^{d}}\tilde{f}_{s+1}(\mathbf{w})\mathrm{d}\mathbf{w}\leq+\infty$. Moreover, we denote the distribution whose density function is $\tilde{f}_{s+1}(\mathbf{w})/\tilde{Z}_{s+1}$ as $\tilde{P}_{s+1}$. Then, we can rewrite the equation (F.13) as follows:
@@
By plugging (F.14) into (F.10), we have
@@
In the above, we can bound the last term, $-\log\tilde{Z}_{s+1}$, by
@@
where $\widehat{P}_{s+1}=\mathcal{N}(\mathbf{w}_{s+1},cH_{s}^{-1})$ and $\widehat{Z}_{s+1}=\int_{\mathbb{R}^{d}}\exp\left(-\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}\right)\mathrm{d}\mathbf{w}$. In (F.16), the inequality holds due to Jensen’s inequality, and the last inequality is by the fact that $\widehat{Z}_{s+1}=\int_{\mathbb{R}^{d}}\exp\left(-\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}\right)\mathrm{d}\mathbf{w}=\sqrt{(2\pi)^{d}c|H_{s}^{-1}|}=Z_{s}$.
By applying the Cauchy-Schwarz inequality, we can further bound the second term on the right-hand side of (F.16) by
@@
and $\lambda_{j}\left(H_{s}^{-1}\right)$ denotes the $j$-th largest eigenvalue of $H_{s}^{-1}$. Then, we can bound the term (a)-1 in (F.17) as follows:
@@
where the inequality holds due to $\mathbb{E}_{X_{j},X_{j^{\prime}}\sim\mathcal{N}(0,1)}[X_{j}^{2}X_{j^{\prime}}^{2}]\leq 3$ for all $j,j^{\prime}\in[d]$, and the last equality holds because $\sum_{j}^{d}\bar{\lambda}_{j}=\operatorname{Tr}\left(c\bar{H}_{s}^{-1}\right)$. Here, $\operatorname{Tr}(A)$ denotes the trace of the matrix $A$.
@@
By plugging (F.19) and (F.20) into (F.17), we have
@@
Combining (F.15), (F.16), and (F.21), and taking summation over $s$, we derive that
@@
$=\sum_{s=1}^{t}\ell_{s}(\mathbf{w}_{s+1})+\frac{1}{2c}\sum_{s=1}^{t}\|\mathbf{w}_{s}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}+\sqrt{6}c\log\left(\frac{\operatorname{det}(M_{t+1,h})}{\operatorname{det}\left(\frac{\lambda}{2}\mathbf{I}_{d}\right)}\right)$

### v9

@@ F.1 Main Proof of Lemma 1
where the first inequality holds since $H_{s}\succeq\lambda\mathbf{I}_{d}$.
By setting $\eta=\frac{1}{2}\log(K+1)+2$ and $\lambda=84\sqrt{2}d\eta$, we derive that
$\|\mathbf{w}_{t}-\mathbf{w}^{\star}\|_{H_{t}}$ $\leq\|\mathbf{w}_{t}-\mathbf{w}_{t+1}\|_{H_{t}}+\|\mathbf{w}_{t+1}-\mathbf{w}^{\star}\|_{H_{t}}$
$\leq\|\mathbf{w}_{t}-\mathbf{w}_{t+1}\|_{H_{t}}+\|\mathbf{w}_{t+1}-\mathbf{w}^{\star}\|_{H_{t+1}}$
$\leq\|\mathbf{w}_{t}-\mathbf{w}_{t+1}\|_{H_{t}}+\beta^{\prime}_{t+1}(\delta)$
$\leq\frac{2\eta}{\lambda}+\beta^{\prime}_{t+1}(\delta):=\beta_{t}(\delta)=\mathcal{O}\left(\sqrt{d}\log t\log K\right).$
where the third inequality follows from the definition of $\beta^{\prime}_{t+1}(\delta)$ in Equation (F.4), and the last inequality holds by Lemma F.9.
This concludes the proof of Lemma 1. ∎
Corollary F.1.
If the boundedness assumption on the parameter is relaxed to $\|\mathbf{w}\|_{2}\leq B$, then $\beta_{t}(\delta)=\mathcal{O}\left(B\sqrt{d}\log t\log K+B^{3/2}\sqrt{d\log K}\right)$.
Proof of Corollary F.1.
When $\|\mathbf{w}\|_{2}\leq B$, following the same analysis as in Lemma 1, we can set $\eta=\frac{1}{2}\log(K+1)+B+1$, $c=7\eta/6$, and $\lambda=\mathcal{O}(dB\log K)$. Under these settings, we have
$\|\mathbf{w}_{t+1}-\mathbf{w}^{\star}\|_{H_{t+1}}$
$\leq\sqrt{2\eta\Bigg[\!11\cdot\left(3\log(1+(K+1)t)+B+2\right)\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+2+\sqrt{6}cd\log\left(1+\frac{t+1}{2\lambda}\right)\!\Bigg]+\!4B^{2}\lambda\!}$
$=\mathcal{O}\left(B\sqrt{d}\log t\log K+B^{3/2}\sqrt{d\log K}\right).$
which concludes the proof. ∎
@@ F.2 Proofs of Lemmas for Lemma 1
Let $\tilde{\ell}_{s}(\mathbf{w})=\ell_{s}(\mathbf{w}_{s})+\langle\nabla\ell_{s}(\mathbf{w}_{s}),\mathbf{w}-\mathbf{w}_{s}\rangle+\frac{1}{2}\|\mathbf{w}-\mathbf{w}_{s}\|^{2}_{\nabla^{2}\ell_{s}(\mathbf{w}_{s})}$ be a second-order approximation of the original function $\ell_{s}(\mathbf{w})$ at the point $\mathbf{w}_{s}$. The update rule in Equation (3) can also be expressed as
@@
$\langle\nabla\tilde{\ell}_{s}(\mathbf{w}_{s+1}),\mathbf{w}_{s+1}-\mathbf{w}^{\star}\rangle\leq\frac{1}{2\eta}\left(\|\mathbf{w}_{s}-\mathbf{w}^{\star}\|_{H_{s}}^{2}-\|\mathbf{w}_{s+1}-\mathbf{w}^{\star}\|_{H_{s}}^{2}-\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\|_{H_{s}}^{2}\right).$ (F.5)
@@
where $\zeta=\log(K+1)+4$. Then, by combining Equation (F.5) and (F.6), we have
@@
Then, one can easily verify that $\|\mathbf{d}^{\prime}_{s}\|_{\infty}\leq 1$ since $\|\mathbf{z}_{s}^{\star}\|_{\infty}\leq\max_{i\in S_{s}}\|x_{si}\|_{2}\|\mathbf{w}^{\star}\|_{2}\leq 1$ and $\|\tilde{\mathbf{z}}_{s}^{\mu}\|_{\infty}\leq\log(1+(K+1)/\mu)$. On the other hand, $\mathbf{d}_{s}^{\prime}$ is $\mathcal{F}_{s}$-measurable since $\mathbf{z}_{s}^{\star}$ and $\tilde{\mathbf{z}}_{s}^{\mu}$ are independent of $\mathbf{y}_{s}$. Moreover, we have $\|\mathbf{d}^{\prime}_{s}\|^{2}_{\mathbb{E}[\bm{\varepsilon}_{s}\bm{\varepsilon}_{s}^{\top}\mid\mathcal{F}_{s}]}=\|\mathbf{d}_{s}\|^{2}_{\mathbb{E}\left[(\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s})(\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s})^{\top}\mid\mathcal{F}_{s}\right]}=\|\mathbf{d}_{s}\|^{2}_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}$ and $\|\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s}\|_{1}\leq 2$. Thus, by Lemma F.5, with probability at least $1-\delta$, for any $t\geq 1$, we have
@@
$\leq(c_{\mu}+1)\sqrt{\tilde{\lambda}+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}\left(\frac{\sqrt{\tilde{\lambda}}}{4}+\frac{4}{\sqrt{\tilde{\lambda}}}\log\cdot\left(\frac{2\sqrt{1+\frac{1}{\tilde{\lambda}}\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}}{\delta}\right)\right)$
$\leq(c_{\mu}+1)\sqrt{\tilde{\lambda}+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}\cdot\left(\frac{\sqrt{\tilde{\lambda}}}{4}+\frac{4}{\sqrt{\tilde{\lambda}}}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right),$ (F.9)
where in the second inequality, we set $\tilde{\lambda}>1$, and the last inequality holds because $\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}=\mathbf{d}_{s}^{\top}\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})\mathbf{d}_{s}\leq 2$. Then, combining Equation (F.8) and (F.9), we obtain
@@
$\leq(c_{\mu}+1)\sqrt{\tilde{\lambda}+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}\cdot\left(\frac{\sqrt{\tilde{\lambda}}}{4}+\frac{4}{\sqrt{\tilde{\lambda}}}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)-\sum_{s=1}^{t}\frac{1}{c_{\mu}}\|\mathbf{z}_{s}^{\star}-\tilde{\mathbf{z}}_{s}^{\mu}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}$
$\leq(c_{\mu}+1)\sqrt{\tilde{\lambda}+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}\cdot\left(\frac{\sqrt{\tilde{\lambda}}}{4}+\frac{4}{\sqrt{\tilde{\lambda}}}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)-(c_{\mu}+1)\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}$
$\leq(c_{\mu}+1)\left(\tilde{\lambda}+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}\right)+(c_{\mu}+1)\left(\frac{\sqrt{\tilde{\lambda}}}{4}+\frac{4}{\sqrt{\tilde{\lambda}}}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}$
@@
$=(c_{\mu}+1)\left(\frac{17}{16}\tilde{\lambda}+2\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+\frac{16}{\tilde{\lambda}}\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right),$ (F.10)
where the third inequality holds due to the AM-GM inequality. Finally, combining Equation (F.7) and (F.10), by setting $\mu=1/t$ and $\tilde{\lambda}=\frac{16}{\sqrt{17}}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)$, we have
$\sum_{s=1}^{t}\left(\ell_{s}(\mathbf{w}^{\star})-\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})\right)$ $\leq(c_{\mu}+1)(2\sqrt{17}+2)\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+2\mu t$
$\leq 11\cdot\left(3\log(1+(K+1)t)+3\right)\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+2$
@@
$L_{s}(\mathbf{w})\leq\tilde{L}_{s}(\mathbf{w})+e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{2}.$ (F.13)
We define the function $\tilde{f}_{s+1}:\mathcal{W}\rightarrow\mathbb{R}$ as
@@
Then, we can establish a lower bound for the expectation in Equation (F.11) as follows:
@@
where the first inequality holds by Equation (F.13) and the last equality holds by Equation (F.12). We define $\tilde{Z}_{s+1}=\int_{\mathbb{R}^{d}}\tilde{f}_{s+1}(\mathbf{w})\mathrm{d}\mathbf{w}\leq+\infty$. Moreover, we denote the distribution whose density function is $\tilde{f}_{s+1}(\mathbf{w})/\tilde{Z}_{s+1}$ as $\tilde{P}_{s+1}$. Then, we can rewrite Equation (F.14) as follows:
@@
By plugging Equation (F.15) into Equation (F.11), we have
@@
In the above, we can bound the last term, $\log\tilde{Z}_{s+1}$, by
@@
where $\widehat{P}_{s+1}=\mathcal{N}(\mathbf{w}_{s+1},cH_{s}^{-1})$ and $\widehat{Z}_{s+1}=\int_{\mathbb{R}^{d}}\exp\left(-\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}\right)\mathrm{d}\mathbf{w}$. In Equation (F.17), the inequality holds due to Jensen’s inequality, and the last inequality is by the fact that $\widehat{Z}_{s+1}=\int_{\mathbb{R}^{d}}\exp\left(-\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}\right)\mathrm{d}\mathbf{w}=\sqrt{(2\pi)^{d}c|H_{s}^{-1}|}=Z_{s}$.
By applying the Cauchy-Schwarz inequality, we can further bound the second term on the right-hand side of Equation (F.17) by
@@
and $\lambda_{j}\left(H_{s}^{-1}\right)$ denotes the $j$-th largest eigenvalue of $H_{s}^{-1}$. Then, we can bound the term (a)-1 in Equation (F.18) as follows:
@@
where the inequality holds due to $\mathbb{E}_{X_{j},X_{j^{\prime}}\sim\mathcal{N}(0,1)}[X_{j}^{2}X_{j^{\prime}}^{2}]\leq 3$ for all $j,j^{\prime}\in[d]$, and the last equality holds because $\sum_{j=1}^{d}\bar{\lambda}_{j}=\operatorname{Tr}\left(c\bar{H}_{s}^{-1}\right)$. Here, $\operatorname{Tr}(A)$ denotes the trace of the matrix $A$.
@@
By plugging Equation (F.20) and (F.21) into Equation (F.18), we have
@@
Combining Equation (F.16), (F.17), and (F.22), and taking summation over $s$, we derive that
@@
$=\sum_{s=1}^{t}\ell_{s}(\mathbf{w}_{s+1})+\frac{1}{2c}\sum_{s=1}^{t}\|\mathbf{w}_{s}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}+\sqrt{6}c\log\left(\frac{\operatorname{det}(M_{s+1})}{\operatorname{det}\left(\frac{\lambda}{2}\mathbf{I}_{d}\right)}\right)$